OpenStudy (aravindg):

a^2+b^@+c^2=1 then bc+ca+ab lies in interval??

OpenStudy (anonymous):

b^???

OpenStudy (aravindg):

2

OpenStudy (aravindg):

@dumbcow , @experimentX , @dpaInc

OpenStudy (aravindg):

@Mani_Jha

OpenStudy (anonymous):

Use QM>= AM >= GM

OpenStudy (aravindg):

hw??? @Aron_West

OpenStudy (anonymous):

(a+b+c)^2=a2+b2+c2+2(bc+ab+ca) (a+b+c)^2=1+2(bc+ab+ca) (bc+ab+ca)=0.5[(a+b+c)^2 - 1]=0.5(a+b+c)^2 -0.5 then ,what to do????

OpenStudy (anonymous):

Quadratic Mean >= Arithmetic Mean >= Geometric mean

OpenStudy (anonymous):

Also use Harmonic Mean.

OpenStudy (anonymous):

where did you get this q from @aravindG

OpenStudy (experimentx):

must be less than 3 and greater than -3 .. thats for sure.

OpenStudy (anonymous):

how @experimentX???

OpenStudy (experimentx):

a,b,c must be less than 1 and greater than 1

OpenStudy (anonymous):

ok...

OpenStudy (experimentx):

but the problem states that what would be the least upper bound and greatest lower bound.

OpenStudy (aravindg):

....i m not getting dat

OpenStudy (experimentx):

you know the answer???

OpenStudy (aravindg):

no

OpenStudy (aravindg):

@Sarkar it came in my test!!

OpenStudy (experimentx):

Oh .. i found answer on google

OpenStudy (anonymous):

you could use the method of lagrange multipliers

OpenStudy (experimentx):

if you still want to solve then beware of google!!!

OpenStudy (anonymous):

It lies between -1/2 and 1.Simple!

OpenStudy (mani_jha):

\[(a+b+c)^{2}=a ^{2}+b ^{2}+c ^{2}+2(ab+bc+ca)\] Let\[ab+bc+ca=x\] Thus RHS becomes 1+2x AM>=GM \[(1+2x)/2\ge \sqrt{2x}\] Now square, factorize and obtain a range. Thanks to @Aron_West and @Sarkar for the hints.

OpenStudy (anonymous):

cool @mani...

OpenStudy (anonymous):

The only pre-requites are ": Any square of a real no.>0 & AM>= GM

OpenStudy (aravindg):

thx all

OpenStudy (anonymous):

@Mani_Jha nice one :D

OpenStudy (anonymous):

To get the bounds : (a+b+c)^2 >=0 or,1+2x>=0 or x>=-1/2 Now use AM>=GM

OpenStudy (mani_jha):

Thanks @anonymoustwo44, and @Aron_West, yes even that condition is to be considered.